Two bicyclists who live 10 miles apart on the same road travel the exact same route every Sunday morning. Roberta's speed is 11 mph and Jeff's speed is 15 mph. They both leave at 8 AM and they meet at the same location for a cup of coffee. How far have they biked and what time is it when they meet for coffee?
Answer by checkley77(12844)

D=RT
10=(11+15)T
10=26T
T=10/26
T=.3846 HOURS IS THE TRAVEL TIME.
11*.3846=4.238 MILES IS THE DISTANCE FOR THE SLOWER BIKER.
15*.3846=5.769 MILES IS THE DISTANCE FOR THE FASTER BIKER.
11:00+.3846*60=11:00+23.076=11:23 THEY WILL MEET.